OpenStudy (anonymous):

Part A: A graph passes through the points (0, 4), (1, 8), and (2, 10). Does this graph represent a linear function or a non-linear function? Explain your answer in words. Part B: Write one example of a linear function and one example of a non-linear function. (Use x and y as the variables)

OpenStudy (anonymous):

someone please help

OpenStudy (anonymous):

well linear you can think line (which have the same slope).... so you need to check to see if those points have the same slope. do you know how to find slope?

OpenStudy (anonymous):

@zonazoo not really

OpenStudy (anonymous):

well points have coordinates (x,y) the formula for slope between any two points of (x, y) is \[\frac{ y _{2}-y _{1} }{ x _{2}-x _{1} }\]

OpenStudy (anonymous):

so pick any two of the three points to test... take the first two (0,4) and (1,8) well 0 is the x1 (first x value) 1 is the x2 (second x value) 4 is the y1 (first y value) and 8 is the y2 (second y value) so plug them into the formula \[\frac{ 8-4 }{ 1-0 }\]

OpenStudy (anonymous):

and we get 4/1 = 4... so the slope between those two points is 4.... now do the same thing with the 2nd and 3rd point and find that slope... if it is the same number then the function is linear, if not... then its not linear.

OpenStudy (anonymous):

I got -1 and -2

OpenStudy (anonymous):

@Zonazoo

OpenStudy (anonymous):

someone please help ill give a medal

OpenStudy (superdavesuper):

-2/-1 = ?